# ref.bib UUITP-22/24

### v1

@@ B.2 Monodromy of the quantum Lebesgue measure
$y_{j}(\mathrm{e}^{-(c\pm\pi\mathrm{i})})=\mathrm{e}^{\mp\pi\mathrm{i}\alpha_{j}}y_{j}(-\mathrm{e}^{-c})\,.$ (B.13)
Applying this result systematically to every term in the expansion of $\mathcal{H}^{D}$, we conclude that
$\Delta_{c}\left[\mathcal{H}^{D}\right]_{0}=\frac{1}{A_{2}\cdot\epsilon}\left(A_{1}\cdot\epsilon-\sum_{i}\epsilon_{i}\sum_{j=1}^{k}\alpha_{j}\right)\sum_{j=1}^{k}\log(-y_{j}(-x))+\sum_{j=1}^{k}\alpha_{j}\log(-y_{j}(-x))+\dots$ (B.14)
@@
Appendix C Quantum volumes as equivariant $B$-model integrals
In this appendix, we collect certain expressions for the three different types of quantum equivariant volumes that play a central role in our work, namely $\mathcal{F}^{D},\mathcal{H}^{D}$ and $\mathcal{K}^{D}$. We postpone an interpretation of these identities to a separate publication.
C.1 Symplectic quotient operators
Let us first recall the Euler integral representation
$\operatorname{\Gamma}(z)=\int_{0}^{\infty}\mathrm{e}^{-u}u^{z-1}\mathop{}\!\mathrm{d}u\,.$ (C.1)
$f(\widehat{u})\cdot\operatorname{\Gamma}(z)=\int_{0}^{\infty}\mathrm{e}^{-u}u^{z-1}f(u)\mathop{}\!\mathrm{d}u\,,$ (C.2)
for any algebraic function $f(u)$.
$\mathcal{G}_{q}(t)\cdot\Psi(\epsilon_{1},\dots,\epsilon_{n}):=$ $\,\delta\Big{(}t+\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial\epsilon_{i}}\Big{)}\cdot\Psi(\epsilon_{1},\dots,\epsilon_{n})$ (C.3)
$\mathcal{G}_{q}(t)\cdot\prod_{i}\operatorname{\Gamma}(\epsilon_{i})=\delta\Big{(}\log\Big{(}\mathrm{e}^{t}\prod_{i}\widehat{u}_{i}{}^{q_{i}}\Big{)}\Big{)}\cdot\prod_{i}\operatorname{\Gamma}(\epsilon_{i})$ (C.4)
$\mathcal{G}_{q}(t)^{-1}\cdot f(t,\epsilon):=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}t\,f(t,\epsilon)\,.$ (C.5)
C.2 Relations among quantum equivariant volumes
$\mathcal{F}^{D}(t,\epsilon)=\Big{(}\prod_{a=1}^{r}\mathcal{G}_{Q^{a}}(t^{a})\Big{)}\cdot\prod_{i=1}^{r+3}\operatorname{\Gamma}(\epsilon_{i})$ (C.6)
$\mathcal{H}^{D}(t,c_{\alpha},\epsilon)=\mathcal{G}_{q^{\alpha}}(c_{\alpha})\cdot\mathcal{F}^{D}(t,\epsilon)\,.$ (C.7)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon):=\Big{(}\prod_{\alpha=1,2}\mathcal{G}_{q^{\alpha}}(c_{\alpha})\Big{)}\cdot\mathcal{F}^{D}(t,\epsilon)$ (C.8)
These relations can be inverted by integrating over the moment map moduli using the operator (C.5). We find
$\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{2}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$ $=\mathcal{H}^{D}(t,c_{1},\epsilon)$ (C.9)
C.3 Equivariant $B$-model integral for $\mathcal{K}^{D}$
We now give explicit integral formulae for the quantum equivariant volumes of the CY3 and of its cuts. These fit into the hierarchy given in (C.9), with the most fundamental object being $\mathcal{K}^{D}$, which will be out starting point.
We begin by rewriting (4.11) by means of the integral representation (C.1) of Euler’s $\operatorname{\Gamma}$ function and in terms of symplectic quotient operators (C.3)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)=\prod_{\alpha=1,2}\int_{\mathrm{i}\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\varphi_{\alpha}}{2\pi\mathrm{i}}\mathrm{e}^{c_{\alpha}\varphi_{\alpha}}\prod_{a=1}^{r}\int_{\mathrm{i}\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\phi_{a}}{2\pi\mathrm{i}}\mathrm{e}^{t^{a}\phi_{a}}\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}+\sum_{a}Q^{a}_{i}\,\phi_{a}+\sum_{\alpha}q^{\alpha}_{i}\,\varphi_{\alpha}}\,\mathrm{e}^{-u_{i}}\\ =\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}}\,\mathrm{e}^{-u_{i}}\,\prod_{a=1}^{r}\delta\Big{(}t^{a}+\sum_{i}Q^{a}_{i}\log u_{i}\Big{)}\prod_{\alpha=1,2}\delta\Big{(}c_{\alpha}+\sum_{i}q^{\alpha}_{i}\log u_{i}\Big{)}\,.$ (C.10)
$\mathcal{K}^{D}(\widetilde{t},\epsilon)$ $=\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}}\,\mathrm{e}^{-u_{i}}\,\prod_{a=1}^{r+2}\delta\Big{(}\widetilde{t}^{a}+\sum_{i}\widetilde{Q}^{a}_{i}\log u_{i}\Big{)}\,.$ (C.11)
$\log u_{i}=-\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}(\widetilde{t}^{a}+\widetilde{Q}^{a}_{j}\log u_{j})\hskip 30.0pt\forall i\neq j\,,$ (C.12)
$u_{i}=\prod_{a=1}^{r+2}\left(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}}\right)^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}\,.$ (C.13)
This allows us to reduce the integral to a single variable
$\mathcal{K}^{D}(\widetilde{t},\epsilon)=\int_{0}^{\infty}\,\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}^{1-\epsilon_{j}}}\left(\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\sum_{i\neq j}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\right)\mathrm{e}^{-u_{j}-\sum_{i\neq j}\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}}\\ =\mathrm{e}^{-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{t}^{a}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\epsilon_{j}-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{Q}^{a}_{j}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\mathrm{e}^{-u_{j}-\sum_{i\neq j}\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}}\,.$ (C.14)
  • $[R(j)]^{i}_{k}=\delta^{i}_{k}$, where $R(j)$ is the square matrix obtained from $R$ by removing the $j$-th column. This trivially follows by noting that $R(j)=\widetilde{Q}(j)^{-1}\widetilde{Q}(j)=\mathbb{1}$.
    $R^{i}_{j}=\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\widetilde{Q}_{j}^{a}=-\sum_{k\neq j}\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\widetilde{Q}_{k}^{a}=-\sum_{k\neq j}\delta^{i}_{k}=-1$ (C.15)
Using these facts, the above expression simplifies to
$\mathcal{K}^{D}(\widetilde{t},\epsilon)$ $=\mathrm{e}^{-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{t}^{a}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\sum_{i=1}^{r+3}\epsilon_{i}}\,\exp\left(-u_{j}\left(1+\sum_{i\neq j}\prod_{a=1}^{r+2}\widetilde{z}_{a}^{\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}\right)\right)$ (C.16)
Let us define
$H(x_{1},x_{2},z):=1+\sum_{i\neq j}\prod_{a=1}^{r+2}\widetilde{z}_{a}^{\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}$ (C.17)
and
$A_{\alpha}^{i}:=\left\{\begin{array}[]{ll}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{r+\alpha}&\text{if $i\neq j$}\\ 0&\text{if $i=j$}\end{array}\right.\,,\hskip 30.0ptM_{a}^{i}:=\left\{\begin{array}[]{ll}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}&\text{if $i\neq j$, $a=1,\dots,r$}\\ 0&\text{if $i=j$}\end{array}\right.$ (C.18)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$ $=\mathrm{e}^{-t\cdot M\cdot\epsilon}x_{1}^{A_{1}\cdot\epsilon}x_{2}^{A_{2}\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\sum_{i=1}^{r+3}\epsilon_{i}}\mathrm{e}^{-u_{j}H(x_{1},x_{2},z)}$ (C.19)
C.4 $H(x_{1},x_{2})$ vs the Hori–Vafa mirror curve
The polynomial $H(x_{1},x_{2})$ appearing in the formula for $\mathcal{K}^{D}$ is related to, but does not quite coincide with, the Hori–Vafa mirror curve. More precisely, since $\mathcal{K}^{D}$ is defined by a double cut, which corresponds a pair of toric Lagrangians, there are two mirror curves that describe the moduli spaces of branes $L_{1}$ and $L_{2}$ associated to the cut.
First of all, the branes may come with independent framings and different choices of coordinates. Therefore their moduli spaces would be given by different expressions $F_{1}(x_{1},y_{1})=0$ and $F_{2}(x_{2},y_{2})=0$. Neither of these functions is equal to $H(x,y)$ on general grounds. However, all three polynomials are related, since they all come from resolving the moment map constraints, on the universal curve $\sum_{i=1}^{r+3}u_{i}=0$. We illustrate this point with an explicit computation.
An example.
$\left[\begin{array}[]{ccc|c}z_{1}&z_{2}&z_{3}&\\ \hline\cr 0&1&-1&c_{1}\\ 1&-1&0&c_{2}\\ \end{array}\right]$ (C.20)
$u_{i}=\exp(-p_{i}+\mathrm{i}\theta_{i})\,,\hskip 30.0pti=1,2,3$ (C.21)
and use the moment map equations
$p_{2}-p_{3}=c_{1}\,,\hskip 30.0ptp_{1}-p_{2}=c_{2}\,,$ (C.22)
to identify the open string moduli as follows
$x_{1}:=\mathrm{e}^{-c_{1}}=\frac{u_{2}}{u_{3}}\,,\hskip 30.0ptx_{2}:=\mathrm{e}^{-c_{2}}=\frac{u_{1}}{u_{2}}\,.$ (C.23)
$y_{1}=\frac{u_{1}}{u_{3}}\,,\hskip 30.0pty_{2}=\frac{u_{3}}{u_{2}}\,.$ (C.24)
The mirror curves for each brane are obtained by Hori–Vafa as follows
$\begin{split}u_{1}+u_{2}+u_{3}&=u_{3}(1+y_{1}+x_{1})=:u_{3}F_{1}(x_{1},y_{1})\\ &=u_{2}(1+y_{2}+x_{2})=:u_{2}F_{2}(x_{2},y_{2})\\ \end{split}$ (C.25)
$x_{2}=y_{1}x_{1}^{-1}\,,\hskip 30.0pty_{2}=x_{1}^{-1}$ (C.26)
$\begin{split}&p_{2}-p_{3}=-\log u_{2}+\log u_{3}=c_{1}\quad\Rightarrow\quad u_{2}=u_{3}\,\mathrm{e}^{-c_{1}}=u_{3}x_{2}\\ &p_{1}-p_{2}=-\log u_{1}+\log u_{2}=c_{2}\quad\Rightarrow\quad u_{1}=u_{2}\,\mathrm{e}^{-c_{2}}=u_{3}x_{1}x_{2}\,.\end{split}$ (C.27)
The Hori–Vafa universal curve can be written as
$u_{1}+u_{2}+u_{3}=u_{3}(x_{1}x_{2}+x_{2}+1)=:u_{3}H(x_{1},x_{2})$ (C.28)
$H(x_{1},x_{2})=F_{1}(x_{1},y_{1})=y_{2}^{-1}F_{2}(x_{2},y_{2})\,.$ (C.29)
C.5 Equivariant $B$-model integrals for $\mathcal{H}^{D}$ and $\mathcal{F}^{D}$
From (C.19), we can immediately obtain $\mathcal{H}^{D}$ and $\mathcal{F}^{D}$ by integration (C.9)
$\mathcal{H}^{D}(t,c_{1},\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x_{1}^{A_{1}\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{2}}{x_{2}}\frac{x_{2}^{A_{2}\cdot\epsilon}}{H(x_{1},x_{2},z)^{\sum_{i}\epsilon_{i}}}\,,$ (C.30)
and
$\mathcal{F}^{D}(t,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{1}}{x_{1}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{2}}{x_{2}}\frac{x_{1}^{A_{1}\cdot\epsilon}x_{2}^{A_{2}\cdot\epsilon}}{H(x_{1},x_{2},z)^{\sum_{i}\epsilon_{i}}}\,.$ (C.31)
From the discussion of Appendix C.4, it follows that we can replace $H(x,y,z)$ by the mirror curve $F(x,y)$ of either brane defining the double cut (or of the brane defining the single cut associated to $\mathcal{H}^{D}$). The formula remains the same, except that $A_{\alpha},M$ get shifted by some factors, see (C.29) for an example.

### v2

@@ C.2 Monodromy of the quantum Lebesgue measure
$y_{j}(\mathrm{e}^{-(c\pm\pi\mathrm{i})})=\mathrm{e}^{\mp\pi\mathrm{i}\alpha_{j}}y_{j}(-\mathrm{e}^{-c})\,.$ (C.13)
Applying this result systematically to every term in the expansion of $\mathcal{H}^{D}$, we find that the only regular instantonic contributions to the monodromy come from terms of the form
$\Delta_{c}\left(\left[\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\right]_{n_{1}}\times\left[\frac{\operatorname{\Gamma}(\sum_{i}\epsilon_{i})\operatorname{\Gamma}(A_{2}\cdot\epsilon)\operatorname{\Gamma}(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon)}{\operatorname{\Gamma}(k\sum_{i}\epsilon_{i})}\right]_{-2}\times\left[F^{(k)}_{D}\right]_{n_{2}}\right)$ (C.14)
where $n_{1},n_{2}$ are positive integers such that $n_{1}+n_{2}=2$, and the middle term in the product is just a constant on which $\Delta_{c}$ acts trivially.
From an explicit computation of the monodromy of these three terms, we find
$\Delta_{c}\left(\left[\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\right]_{2}\times\left[F^{(k)}_{D}\right]_{0}\right)=\sum_{i}\epsilon_{i}\Big{(}A_{1}\cdot\epsilon-\sum_{i}\epsilon_{i}\sum_{j=1}^{k}\alpha_{j}\Big{)}\sum_{j=1}^{k}\log(-y_{j}(-x))+\dots$ (C.15)
$\Delta_{c}\left(\left[\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\right]_{1}\times\left[F^{(k)}_{D}\right]_{1}\right)=-\frac{A_{2}\cdot\epsilon}{k}\Big{(}A_{1}\cdot\epsilon-2\sum_{i}\epsilon_{i}\sum_{j=1}^{k}\alpha_{j}\Big{)}\sum_{j=1}^{k}\log(-y_{j}(-x))+\dots$ (C.16)
[…]
Summing them all together and multiplying by the constant, we finally conclude that
$\Delta_{c}\left[\mathcal{H}^{D}\right]_{0}=\frac{1}{A_{2}\cdot\epsilon}\left(A_{1}\cdot\epsilon-\sum_{i}\epsilon_{i}\sum_{j=1}^{k}\alpha_{j}\right)\sum_{j=1}^{k}\log(-y_{j}(-x))+\sum_{j=1}^{k}\alpha_{j}\log(-y_{j}(-x))+\dots$ (C.18)
@@
We should remark here that derivation of the formula for the monodromy of $\mathcal{H}^{D}$ makes use of the explicit series expansion (A.1) for the Lauricella function $F_{D}^{(n)}$ in (2.30), whose domain of convergence is the product of unit disks
$|1+y_{j}^{-1}|<1\qquad j=1,\dots,k\,.$ (C.19)
In particular, this condition must hold for the asymptotic values of the roots $y_{j}(x)$ within some region of the large volume regime, as in (2.31).
Whenever condition (C.19) is not satisfied, we find ourselves in a region of parameters space where the series expansion (A.1) is no longer valid. In this case, we need to compute the series expansion of $F_{D}^{(n)}$ in the new region by means of analytic continuation formulas, see e.g. [Bezrodnykh_2018] and references therein. Once an appropriate series expansion has been found, we can then repeat the steps in this Appendix to compute the correct formula for the monodromy compatible with the values of the parameters. In [Cassia:2023uvd], we address one example when this issue arises in the case of local $\mathbb{P}^{2}$, where the equivariant disk potential is computed in a phase in which (C.19) is violated. In that instance, the Lauricella function specializes to a simpler hypergeometric series, whose analytic continuations are well-known.